A four digit number is to be formed from the digits 1, 3, 5 and 8. What is the probability that the number so formed will:
a) start with 5?
b) be an odd number? and
c) be greater than 5000?

Answer by stanbon(75887)

A four digit number is to be formed from the digits 1, 3, 5 and 8. What is the probability that the number so formed will:
a) P(start with 5) = (1*3!)/4! = 1/4
---------------------------------------
b) P(be an odd number) = (3*3*2*1)/4! = 18/24 = 3/4
----------------------------------------------------
c) P(be greater than 5000) = (1*3!)/4! = 1/4
------------------------
Cheers,
Stan H.